OpenStudy (anonymous):

rectangular container with no top is to have a volume of 8m^2. Find the dimensions that will minimize the amount of material used. This is in terms on xyz.

OpenStudy (anonymous):

So this is what i did... xyz=8 z=8/xy A= xy+2xz+2yz then sub in the z

OpenStudy (anonymous):

A(x,y) xy+16/y +16/x I'm having probs with finding the values for x,y, z

OpenStudy (anonymous):

Ax= y-16/x^2 Ay=x-16/y^2

terenzreignz (terenzreignz):

How about considering f(x,y,z) instead? f(x,y,z) = xy + 2xz + 2yz and find fx, fy, and fz ?

OpenStudy (anonymous):

i assume you mean 8 m^3 right?

OpenStudy (anonymous):

in my text it says m^2 but i think its a mistake

OpenStudy (anonymous):

do you know lagrangian multipliers? .. what math class is this?

OpenStudy (anonymous):

nope its 2nd part of calculus 1

OpenStudy (anonymous):

integrals

OpenStudy (anonymous):

so as @terenzreigns said, why not consider a general rectangle with dimensions x,y,z what is the volume of that?

OpenStudy (anonymous):

and the surface area of that (surface area = amount of material used in this case)

OpenStudy (anonymous):

isn't that just A= xy+2xz+2yz ?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

so how can we minimize that while ensuring that x*y*z = 8?

OpenStudy (anonymous):

find the zeros?

OpenStudy (anonymous):

make z= 8/xy then sub them into A= xy+2xz+2yz

OpenStudy (anonymous):

so that's one way you can do it. you can then find a minimum of that function. In[1]:= (x*y + 2*x*z + 2*y*z) /. z -> 8/(x*y) In[2]:= f = 16/x + 16/y + x y Out[2]= 16/x + 16/y + x y In[30]:= NSolve[D[f, x] == 0 && D[f, y] == 0, {x, y}, Reals] Out[30]= {{x -> 2.51984, y -> 2.51984}} then you solve for z = 8/(x*y) = 1.25992 this gives you a minimized surface area of 19.0488